Metallica Bearings, Inc., is a young start-up company. No dividends will be paid on the stock over the next nine years because the firm needs to plow back its earnings to fuel growth. The company will pay a dividend of $15 per share 10 years from today and will increase the dividend by 5 percent per year thereafter. If the required return on this stock is 14.5 percent, what is the current share price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Value at year 9 = D10 / required rate - growth rate

Value at year 9 = 15 / 0.145 - 0.05

Value at year 9 = 15 / 0.095

Value at year 9 = 157.894737

Current share price = FV / (1 + r)n

Current share price = 157.894737 / (1 + 0.145)9

Current share price = 157.894737 / 3.38259

Current share price = $46.68